OpenStudy (faolan):

Help? PLS??? Use the process of completing the square to fill in the blanks. y = x^2 + 8x + 25 y = (x +BLANK)^2 + BLANK

OpenStudy (faolan):

@RhondaSommer

OpenStudy (faolan):

@rEdNeCkGiRl101

OpenStudy (redneckgirl101):

what grade is this

OpenStudy (faolan):

Algebra 2

OpenStudy (faolan):

11

OpenStudy (faolan):

i suck at math

OpenStudy (redneckgirl101):

what the name of the unit

OpenStudy (faolan):

Completing the Square

OpenStudy (redneckgirl101):

hold on

OpenStudy (faolan):

The Unit is Real Numbers. But the lesson is called Completing the Square

OpenStudy (redneckgirl101):

https://mathway.com go to this website

OpenStudy (faolan):

i have to pay to get the answer........

OpenStudy (redneckgirl101):

i dont hold on

OpenStudy (redneckgirl101):

i dont get this lesson so im going to use it

OpenStudy (faolan):

ok. U know what im talking about?

OpenStudy (redneckgirl101):

no

OpenStudy (faolan):

oh

OpenStudy (redneckgirl101):

im so confused

OpenStudy (faiqraees):

\[\large\rm y=(x^2 +8x)+25\]\[\large\rm y=(x^2 +8x +(8/2)^2)+25-(8/2)^2\]\[\large\rm y=(x^2 +8x +(4)^2)+25-(4)^2\] Can you proceed?

OpenStudy (redneckgirl101):

thanks ^ @FaiqRaees

OpenStudy (faolan):

What would I do next?

OpenStudy (faolan):

@FaiqRaees

OpenStudy (faiqraees):

\[\large\rm y=((x+4)^2)+25-16\]\[\large\rm y=(x+4)^2+9\]

OpenStudy (faolan):

thats it?

OpenStudy (faolan):

i feel dumb now. lol. thx for the help!